OpenStudy (anonymous):

8.6X + 4 • (100 - X) = 842

OpenStudy (anonymous):

(8.6)x + 4(100-x) = 842 (8.6)x + 400 - 4x = 842 (4.6)x = 442 x = 442/4.6 Simplify the fraction